Analysis of test beam results

Position 22 masked for all the �les. RP1 mouns20 RP1 muon31 RP2 muon5

RP2 muon6 th10

events total  10550  5385  22788  16623
non-empty events  8220  4469  16949  12583

events with at least 3 active planes per event 5359 3519 14790 11266 events with at least 3 active planes

per event after track selection  5224  3195  14329  10929
selected hits (clusters)  49410  31312  133537  102694
rejected hits (clusters)  6124  3742  13480  10028

Table 1: Events and hits statistics. The counts correspond to the entire run.

Figure 1: The numbers of active planes per event. Left: u projection, right: v projection.

Figure 2: 2D histograms of the numbers of active planes.

Figure 3: The histogram of the number of rejected hits (clusters) per event.